OpenStudy (anikate):

use a property of logarithms to evaluate the expression. ln e^3

OpenStudy (ipwnbunnies):

\[\log(a^{b}) = b*\log(a)\] Also, it's help to know that ln(e) = 1

OpenStudy (ipwnbunnies):

helpful*

OpenStudy (anikate):

oh ok thanks :) so I can just pull out the 3, so it would be 3 * ln(e) = 3*1 = 3

OpenStudy (ipwnbunnies):

Good job man. :)

OpenStudy (anikate):

yup thanks for the help